A solution contains a solute with molar mass 94.5 g/mol. The solution is
1.77 m and also 1.70 M. Calculate the density of the solution.
1
Expert's answer
2012-06-19T09:17:11-0400
c - molarity = 1.70 M
b - molality = 1.77 m
M - molar mass = 94.5
g/mol
d - density - g/mL

c = n(solute) / V(solution)
n(solute) =
c*V(solution)
m(solute) = n(solute)*M(solute)
assume, that we have 1L of
solution. Hence,
n(solute) = 1.7 M * 1 L = 1.7 mol
m(solute) = 1.7 mol *
94.5 g/mol = 160.65 g

b = n(solute) / m(solvent)
m(solvent) =
n(solute) / b
m(solvent) = 1.7 mol / 1.77 = 960.45 g

m(solution) =
m(solute) + m(solvent) = 160.65 g + 960.45 g = 1121.1 g
d(solution) =
m(solution) / V(solution) = 1121.1 g / 1 L = 1.12 kg/L = 1.12 g/mL